# When large language models meet evolutionary algorithms

### v2

@@ Abstract
Pre-trained large language models (LLMs) have powerful capabilities for generating creative natural text. Evolutionary algorithms (EAs) can discover diverse solutions to complex real-world problems. Motivated by the common collective and directionality of text generation and evolution, this paper illustrates the parallels between LLMs and EAs, which includes multiple one-to-one key characteristics: token representation and individual representation, position encoding and fitness shaping, position embedding and selection, Transformers block and reproduction, and model training and parameter adaptation. By examining these parallels, we analyze existing interdisciplinary research, with a specific focus on evolutionary fine-tuning and LLM-enhanced EAs. Drawing from these insights, valuable future directions are presented for advancing the integration of LLMs and EAs, while highlighting key challenges along the way. These parallels not only reveal the evolution mechanism behind LLMs but also facilitate the development of evolved artificial agents that approach or surpass biological organisms.
@@
Figure 1: Both tokens in a text and individuals in a population can be regarded as sequences. In a text, each token corresponds to a specific position, while in a population, each individual is associated with a particular fitness rank. Text sequences have a natural directionality derived from human-defined grammatical rules. Population evolution is also directional, which is primarily driven by individual fitness.
@@
Large language models (LLMs) learn statistical patterns with temporal relations from text sequences in an unsupervised manner to establish probability distributions of texts, such as generative pre-trained Transformer (GPT) [1] and bidirectional encoder representations from transformer (BERT) [2]. Typical LLMs can analyze input tokens and generate the most likely subsequent tokens, a process that iterates to produce creative texts. This sequence-to-sequence model with powerful understanding and generation capabilities has been employed to assist users on a variety of creative tasks, including writing, mathematical discovery, and chemical research [3, 4, 5, 6, 7, 8]. Meanwhile, training LLMs on vast text demands significant computing resources, as exemplified by ChatGPT’s pre-training consumption of several thousand petaflop/s-days [3]. Fine-tuning techniques have been proposed to alleviate the computational challenges of training from scratch [9]. In a model-as-a-service scenario [10], LLMs are only accessed as inference APIs. Thanks to their gradient-free nature, evolutionary algorithms (EAs) are employed to fine-tune LLMs in this black-box scenario, aiming to boost performance on downstream tasks [10]. Evolutionary fine-tuning is renowned for low cost, needing no internal model gradient access.
Figure 2: Overview of the generative pre-training Transformer (GPT) and genetic algorithm (GA). Modules of the same color indicate parallels, as exemplified by the analogy between crossover in GA and attention in GPT. GPT continuously generates subsequent tokens by iterating over a context window. The tokens in the input window are transformed by a large-scale Transformer block (multi-head self-attention and feed-forward neural network). These tokens function collectively, providing contextual information for accurate generation. GA generates individuals with high fitness by iterating a population. Individuals in the parent population are transformed by reproduction operators (crossover and mutation). These individuals exhibit collective intelligence, helping to explore the search space and seek the optimal solutions.
@@
Drawing from biological evolution, EAs continuously maintain evolving systems (population or probability distribution) through reproduction and selection to explore fitness landscapes [11, 12]. Typical methods include the genetic algorithm (GA), evolutionary strategy (ES), and genetic programming (GP) [13, 14]. In principle, only individuals and their fitness are needed to drive the evolutionary process in these approaches. Due to advancements in computational resources, EAs have provided diverse solutions to complex black-box optimization issues, such as neuroevolution [15], industrial design [16, 17, 18, 19], and natural sciences [20]. Nonetheless, most EAs are task-specific, and their capabilities do not automatically increase with experience [21, 22, 23, 24]. Recently, Transformer-enhanced EAs utilize basic Transformer models [25] to learn optimization experiences, while LLM-enhanced EAs employ well-trained LLMs [26, 27] to produce optimization experiences. Both approaches intend to improve the performance and generalization of EAs.
@@
Fig. 1 demonstrates that text and population can be regarded as sequence data, specifically exhibiting directionality. LLMs and EAs are designed to learn or simulate such sequential data. Fig. 2 illustrates the process of the sequences generated by LLMs and EAs, taking the GPT [1] and GA [28] as examples respectively. During the generation process, the context window and population are continuously updated to generate creative texts and diverse solutions. Notably, the context window and population share a collective nature, with constituent elements functioning cohesively in their domains. Inspired by the above directionality and collective, we raise the following issues: Are there parallels between LLMs and EAs? Moreover, current interdisciplinary research, whether focusing on evolutionary fine-tuning or LLM-enhanced EAs, remains in its infancy. Can these parallels guide future interdisciplinary integration? To address these issues, this paper draws analogies between the primary characteristics of LLMs and EAs, emphasizing their common mechanisms. At the micro level, we analyze key interdisciplinary research related to each parallel, which not only supports our analogies but also offers insights for potential improvements. At the macro level, we systematically summarize evolutionary fine-tuning and LLM-enhanced EAs to reveal critical challenges.
@@ 1 Parallels
Table 1: A comparison of large language models and evolutionary algorithms in terms of key characteristics, where ”N/A” indicates a lack of corresponding interdisciplinary research in that area.
@@
Characteristics Classic methods Traits Key interdisciplinary research Characteristics Classic methods Traits Key interdisciplinary research
Token representation One-hot encoding + Linear transformation Collective, Uniqueness, finite [10, 29] Individual representation Real encoding + Random embedding [30] Collective, Uniqueness, Infinite and changing [31, 27]
Position encoding Sine and cosine functions Uniqueness, Relativity N/A Fitness shaping Rank transformation, Utility function [32] Uniqueness, Relativity, Directionality [25]
Position embedding Absolute, Relative, Rotary [33] Relativity N/A Selection Tournament selection, Rank selection Relativity, Directionality [25]
Transformer block Multi-head self-attention + Feed-forward neural network Position-insensitive, Parallelism, Sparsity, token and position information, Singleness, synergy [34] Reproduction Arithmetic crossover + Uniform mutation Fitness-insensitive, Parallelism, Sparsity, Individual and fitness information, Singleness, synergy [25, 35]
Model training Pre-training, Fine-tuning, Reinforcement learning Learning, Exploration, Parameter space, Language space [10, 29] Parameter adaptation Hyper-heuristics, Pre-training, Fine-tuning, Meta-learning Learning, Exploration, Parameter space, Search space [25, 27, 36, 37, 31]
[…]
@@ 1.2 Position encoding and fitness shaping
Position encoding models the dependence of tokens at different positions in the sequence. In GPT [1], sine and cosine functions of different frequencies are adopted to encode dependencies. Each token has a unique position encoding consisting of cosine functions with different frequencies. The combination of several cosine waves contains relative distance information between tokens. However, due to the symmetry of distances, the position encoding cannot distinguish sequence direction.
@@
Both position encoding and fitness shaping share the characteristics of coding uniqueness. Classic position encoding effectively models the relative positions between tokens, although it does not capture the directionality of the sequence. Fitness shaping describes the relative and directional ordering of individual fitness. Inspired by fitness shaping, the integration of sequence directionality into position encoding emerges as a noteworthy research direction. Conversely, in existing Transformer-enhanced EAs [25], fitness values are directly utilized for position encoding within the Transformer model. According to our analogy, the introduction of fitness shaping has the potential to aid Transformer-enhanced EAs in managing selective pressures.
@@ 1.3 Transformer block and reproduction
The multi-head self-attention mechanism operates by combining multiple self-attention mechanisms to focus on information from different subspaces. Many studies show that the learned attention matrix is sparse. To reduce computational complexity, improvement mechanisms have been proposed, such as sparse attention and linear attention [43]. In Transformer blocks, the FFN enhances the expressive ability of each token $x_{i}^{{}^{\prime}}$ by applying a nonlinear function $f{(x_{i}^{{}^{\prime}})}$. Residual connections help alleviate the vanishing gradient problem, enabling deeper feature learning. Layer normalization stabilizes the training process and improves convergence speed.
@@
Comparing (1) with (4), attention and crossover share a similar mathematical representation, exhibiting parallelism and sparsity. Attention does not explicitly model token positions. Similarly, crossover inherently does not consider individual fitness. The attention and selection matrices play analogous roles: one determines token feature combinations, while the other governs parent genetic combinations. The attention matrix is parameterized based on token embeddings, while the selection matrix is heuristically built on individual relationships. Additionally, (1) demonstrates that the input and output of attention occupy distinct latent spaces. However, (4) reveals the population keeps the same representation space across crossover.
Both the FFN and mutation operate independently on each singleton (token or individual) and can be processed in parallel. Existing works [46, 34] show that the removal of FFNs degrades Transformer performance, emphasizing the crucial role of both attention and FFNs. Analogously, the synergistic effect of crossover and mutation results in the super efficiency of EAs [47, 48]. The randomness of mutations contributes to the generation of diverse individuals within a population. FFNs with dropouts show randomness in training but are deterministic in testing. Introducing randomness into FFNs has the potential to enhance the diversity of LLM outputs.
Zhang et al. [34] first analogized Transformer blocks to reproduction, utilizing dynamic local populations in EAs for a stronger Vision Transformer. Crossover and mutation were modeled with attention and FFN for Transformer-enhanced EAs [35]. These efforts confirm our analogy and demonstrate the potential for progress through idea sharing between advanced Transformers and reproduction.
@@ 1.4 Position embedding and selection
The attention matrix is influenced by both token and position relations. The selection matrix is usually designed based on fitness relations. In complex fitness landscapes, additional considerations such as genetic similarity among individuals are factored into the selection matrix. For instance, in multi-modal optimization with multiple global optima [57, 58], individual distances within the search space are employed as a criterion to preserve population diversity during selection. Therefore, the selection matrix consists of individual relations and fitness relations. Essentially, the striking similarity between attention and selection matrices in handling relational information stems from the analogy drawn to tokens and individuals, as well as to positions and fitness. Existing Transformer-enhanced EA merges individual embeddings with fitness embeddings, echoing how token embeddings are combined with positional embeddings [25]. This practice serves as a support for our analogy. Advanced attention mechanisms, such as those incorporating rotational positional embedding [33], have the potential to enhance the performance of this operation in Transformer-enhanced EAs.
@@ 1.5 Model training and parameter adaptation
where $p_{i:N}$ refers to the $i$th ranked individual in a population with $N$ individuals based on fitness. The first term is the mean update, while the second term is the ranking-$N$ update of the covariance matrix. Recently, Transformer-enhanced EAs adaptively updated parameters from optimization experiences on a set of optimization tasks, improving the generalization ability of EAs on new tasks. Common methods include pre-training [25, 27] and meta-learning [36, 37]. In pre-training, optimization experiences for multi-objective optimization [25] consist of the population and their fitness generated by multi-objective EAs on numerous benchmarks. Optimization experiences for GP [27] include incremental changes in files submitted by humans to version control systems like Github. Meta-learning [36, 37] adaptively updates parameters by optimizing the average performance of EAs across a set of tasks. Regrettably, no comprehensive study has compared these two methods within a single framework. Furthermore, well-trained LLMs are directly utilized as reproduction operators with human-like experience. Prompts based on historical populations are constructed to guide LLMs in generating the desired output population [31, 27].
@@
Despite differing implementation strategies, LLMs and EAs converge on a shared fundamental objective: revealing the underlying probability distributions within data, thereby facilitating the learning and exploration of knowledge. In pre-training and supervised fine-tuning, LLMs construct conditional probability distributions through the accurate prediction of tokens, learning vast pre-existing knowledge. Reinforcement learning adjusts the LLM parameters based on the rewards. Evolutionary fine-tuning automatically searches for high-quality prompts to improve output quality. These two paradigms explore new knowledge specific to the target task. Learning and exploration jointly ensure the generative and generalization abilities of LLMs. EAs shape probability distributions based on fitness evaluated via real-time individual-environment interaction. Traditional EAs continuously explore knowledge specific to a single target task. In Transformer/LLM-enhanced EAs, models learn from existing optimization experiences or human-like experiences, endowing EAs with powerful learning capabilities. Additionally, LLM training works in the parameter space, while evolutionary fine-tuning extends to the language space. EAs operate both in the search space (e.g., GA) and the parameter space (e.g., CMAES). The aforementioned analogy provides a reasonable motivation for interdisciplinary research: merging the exceptional learning capabilities of LLMs with the remarkable exploration abilities of EAs can foster advancements in their respective fields.
@@
In (12), the hyperparameter $\mu$ is carefully tuned manually, as it affects the generalization ability of LLMs. The loss trade-off is modeled as a multi-objective optimization problem [61, 62]. Applying advanced multi-objective EAs aids in creating stronger supervised fine-tuning paradigms. Recent studies [54] show that ES has advantages over gradient-based reinforcement learning for long episodes with very many time steps. ES is promising as an alternative to reinforcement learning for training LLMs in multi-turn dialogue systems. Evolutionary fine-tuning typically focuses on addressing a single-task scenario. Incorporating optimization experiences from various fine-tuning contexts into EAs can boost their adaptability.
@@
1.6 Discussion
Although LLMs and EAs developed independently, their parallels are truly remarkable. From a macro perspective, analogies hold promise for the development of artificial agents capable of learning from established knowledge while continuously exploring new knowledge. These parallels have been implicitly mentioned or experimentally demonstrated in interdisciplinary research [10, 29, 31, 35, 25, 27, 36, 37, 34]. However, a unified paradigm with one-to-one key feature correspondence has not emerged. Our analogy aims to provide a roadmap for its realization. In existing efforts to integrate EAs and LLMs, evolutionary fine-tuning in black-box scenarios and LLM-enhanced EAs are receiving increasing attention. Next, this paper provides a comprehensive review of them to identify key challenges.

### v3

@@ Abstract
Pre-trained large language models (LLMs) exhibit powerful capabilities for generating natural text. Evolutionary algorithms (EAs) can discover diverse solutions to complex real-world problems. Motivated by the common collective and directionality of text generation and evolution, this paper first illustrates the conceptual parallels between LLMs and EAs at a micro level, which includes multiple one-to-one key characteristics: token representation and individual representation, position encoding and fitness shaping, position embedding and selection, Transformers block and reproduction, and model training and parameter adaptation. These parallels highlight potential opportunities for technical advancements in both LLMs and EAs. Subsequently, we analyze existing interdisciplinary research from a macro perspective to uncover critical challenges, with a particular focus on evolutionary fine-tuning and LLM-enhanced EAs. These analyses not only provide insights into the evolutionary mechanisms behind LLMs but also offer potential directions for enhancing the capabilities of artificial agents.
@@
Figure 1: Both tokens in a text and individuals in a population can be regarded as sequences.
@@
Natural language processing (NLP) focuses on enabling computers to understand, generate, and process human language, covering tasks such as text generation [1], text segmentation [2], named entity recognition [3], sequence labeling [4], and relation extraction [5]. Large language models (LLMs), such as generative pre-trained Transformer (GPT) [6] and bidirectional encoder representations from transformer (BERT) [7], primarily learn statistical patterns with temporal relations from text sequences, often in an unsupervised manner, to establish probability distributions of texts. These models have become foundational tools for the aforementioned NLP tasks. By analyzing input tokens and iteratively generating the most likely subsequent tokens, LLMs like GPT and BERT can produce coherent and contextually relevant texts. This sequence-to-sequence model with powerful understanding and generation capabilities has been employed to assist users on a variety of innovative tasks, including writing, mathematical discovery, and chemical research [8, 9, 10, 11, 12, 13]. Meanwhile, training LLMs on vast text demands significant computing resources, as exemplified by ChatGPT’s pre-training consumption of several thousand petaflop/s-days [8]. Fine-tuning techniques have been proposed to alleviate the computational challenges of training from scratch [14]. In a model-as-a-service scenario [15], LLMs are only accessed as inference APIs. Thanks to their gradient-free nature, evolutionary algorithms (EAs) are employed to fine-tune LLMs in black-box scenarios, where they rely solely on forward propagation and do not require access to internal model gradients [15]. This makes EAs a practical choice for such settings.
Figure 2: Overview of the generative pre-training Transformer (GPT) and genetic algorithm (GA). Modules of the same color indicate parallels, as exemplified by the analogy between crossover in GA and attention in GPT.
@@
Drawing from biological evolution, EAs continuously maintain evolving systems (population or probability distribution) through reproduction and selection to explore fitness landscapes [16, 17, 18]. Typical methods include the genetic algorithm (GA), evolutionary strategy (ES), and genetic programming (GP) [19, 20]. In principle, only individuals and their fitness are needed to drive the evolutionary process in these approaches. Due to advancements in computational resources, EAs have provided diverse solutions to complex black-box optimization issues, such as neuroevolution [21], industrial design [22, 23, 24, 25], and natural sciences [26]. Nonetheless, most EAs are task-specific, and their capabilities do not automatically increase with experience [27, 28, 29, 30]. Recently, Transformer-enhanced EAs utilize basic Transformer models [31] to learn optimization experiences, while LLM-enhanced EAs employ well-trained LLMs [32, 33] to produce optimization experiences. Both approaches intend to improve the performance and generalization of EAs.
Fig. 1 demonstrates that text and population can be regarded as sequence data, specifically exhibiting directionality. In a text, each token corresponds to a specific position, while in a population, each individual is associated with a particular fitness rank. Text sequences have a natural directionality derived from human-defined grammatical rules. Population evolution is also directional, primarily driven by fitness-based selection. LLMs and EAs are designed to learn or simulate such sequential data (text sequence and population sequence). Fig. 2 illustrates the process of the sequences generated by LLMs and EAs, taking the GPT [6] and GA [34] as examples respectively. GPT continuously generates subsequent tokens by iterating over a context window. The tokens in the input window are transformed by a large-scale Transformer block (multi-head self-attention and feed-forward neural network). These tokens function collectively, providing contextual information for accurate generation. GA generates individuals with high fitness by iterating a population. Individuals in the parent population are transformed by reproduction operators (crossover and mutation). These individuals exhibit collective intelligence, helping to explore the search space and seek the optimal solutions. During the generation process, both the context window in GPT and the population in GA are continuously updated to produce coherent texts and diverse solutions, respectively. Notably, the context window and population share a collective nature, where their constituent elements (tokens in GPT and individuals in GA) function cohesively within their respective domains.
@@
Inspired by the above directionality and collective, we raise the following issues: Are there parallels between LLMs and EAs? How can these parallels be utilized to address existing limitations and foster innovation in the coupling of LLMs and EAs? Moreover, since current interdisciplinary research focusing on either evolutionary fine-tuning or LLM-enhanced EAs remains in its infancy, what are the key challenges faced by these efforts? To address these issues, this paper draws conceptual analogies between the primary characteristics of LLMs and EAs, emphasizing their common mechanisms. At the micro level, we analyze key interdisciplinary research related to each parallel, which not only supports our analogies but also offers insights for potential improvements. At the macro level, we systematically summarize evolutionary fine-tuning and LLM-enhanced EAs to reveal critical challenges. The main contributions of this paper are as follows:
  1. 1.
    At the micro level, we analyze existing research by drawing conceptual analogies between the key characteristics of LLMs and EAs, aiming to inspire novel ideas and technologies that can advance these fields.
  2. 2.
    At the macro level, we provide the first systematic overview of evolutionary fine-tuning and LLM-enhanced EAs, highlighting key challenges for future research.
The remainder of this paper is organized as follows. Section 1 presents the conceptual parallels between LLMs and EAs from five perspectives, introducing potential technical improvements. In Sections 2 and 3, we summarize evolutionary fine-tuning in black-box scenarios and LLM-enhanced EAs, respectively. Additionally, key future challenges are identified. Finally, Section 4 summarizes the work in this paper.
@@ 1 Parallels
Table 1: A comparison of large language models and evolutionary algorithms in terms of key characteristics (Chars.), where ”N/A” indicates a lack of corresponding interdisciplinary research in that area.
@@
Sec. Chars. Classic methods Traits Key interdisciplinary research Chars. Classic methods Traits Key interdisciplinary research
1.1 Token representation One-hot encoding + Linear transformation Collective, Uniqueness, finite [15, 35] Individual representation Real encoding + Random embedding [36] Collective, Uniqueness, Infinite and changing [37, 33]
1.2 Position encoding Sine and cosine functions [6] Uniqueness, Relativity N/A Fitness shaping Rank transformation, Utility function [38] Uniqueness, Relativity, Directionality [31]
1.4 Position embedding Absolute, Relative, Rotary [39] Relativity N/A Selection Tournament selection, Rank selection Relativity, Directionality [31]
1.3 Transformer block Multi-head self-attention + Feed-forward neural network Position-insensitive, Parallelism, Sparsity, token and position information, Singleness, synergy [40] Reproduction Arithmetic crossover + Uniform mutation Fitness-insensitive, Parallelism, Sparsity, Individual and fitness information, Singleness, synergy [31, 41]
1.5 Model training Pre-training, Fine-tuning, Reinforcement learning Learning, Exploration, Parameter space, Language space [15, 35] Parameter adaptation Hyper-heuristics, Pre-training, Fine-tuning, Meta-learning Learning, Exploration, Parameter space, Search space [31, 33, 42, 43, 37]
@@ 1.2 Position encoding and fitness shaping
Position encoding models the dependence of tokens at different positions in the sequence. In GPT [6], sine and cosine functions of different frequencies are adopted to encode dependencies. Each token has a unique position encoding consisting of cosine functions with different frequencies. The combination of several cosine waves contains relative distance information between tokens. However, due to the symmetry of distances, the position encoding cannot distinguish sequence direction. Notably, Lyu et al. [48] investigate whether models can learn directionality, emphasizing its importance for interpretability and performance.
@@
Both position encoding and fitness shaping share the characteristics of coding uniqueness. The position encoding used in GPT effectively models the relative positions between tokens, although it does not explicitly capture the directionality of the sequence. Fitness shaping describes the relative and directional ordering of individual fitness. Inspired by fitness shaping, the integration of sequence directionality into position encoding emerges as a noteworthy research direction. Drawing from fitness shaping techniques in CMA-ES/NES [49, 38], future work could design position encodings that preserve order invariance, capturing both the relative positions and directionality of tokens in text. Such methods could improve performance in tasks like long-text generation and question answering by ensuring that the model respects the temporal and contextual order of information. Conversely, in existing Transformer-enhanced EAs [31], fitness values are directly utilized for position encoding within the Transformer model. According to our conceptual analogy, the introduction of fitness shaping has the potential to aid Transformer-enhanced EAs in managing selective pressures.
@@ 1.3 Transformer block and reproduction
The multi-head self-attention mechanism operates by combining multiple self-attention mechanisms to focus on information from different subspaces. Many studies show that the learned attention matrix is sparse [50, 51, 52]. To address the computational cost, various improvement mechanisms have been proposed, such as sparse attention and linear attention [53]. In Transformer blocks, the FFN enhances the expressive ability of each token $x_{i}^{{}^{\prime}}$ by applying a nonlinear function $f{(x_{i}^{{}^{\prime}})}$. Residual connections help alleviate the vanishing gradient problem, enabling deeper feature learning. Layer normalization stabilizes the training process and improves convergence speed.
@@
Comparing (1) with (4), attention and crossover share a similar mathematical representation, exhibiting parallelism and sparsity. Attention does not explicitly model token positions. Similarly, crossover inherently does not consider individual fitness. The attention and selection matrices play analogous roles: one determines token feature combinations, while the other governs parent genetic combinations. The attention matrix is parameterized based on token embeddings, while the selection matrix is heuristically built on individual relationships. Additionally, (1) demonstrates that the input and output of attention occupy distinct latent spaces. However, (4) reveals the population keeps the same representation space across crossover. This analogy highlights opportunities to borrow ideas from attention mechanisms to improve crossover operators and vice versa, as demonstrated in recent studies [41, 40].
Both the FFN and mutation operate independently on each singleton (token or individual) and can be processed in parallel. Existing works [56, 40] show that the removal of FFNs degrades Transformer performance, emphasizing the crucial role of both attention and FFNs. Analogously, the synergistic effect of crossover and mutation results in the super efficiency of EAs [57, 58]. While FFNs and mutation are not mathematically equivalent, their functional roles in maintaining performance are conceptually similar. This analogy suggests potential opportunities to enhance FFNs by introducing controlled randomness, inspired by the role of mutation in generating diversity within populations in EAs.
Zhang et al. [40] first analogized Transformer blocks to reproduction, utilizing dynamic local populations in EAs to enhance Vision Transformers. Similarly, Li et al. [41] modeled crossover and mutation using attention and FFN mechanisms, respectively, to develop Transformer-enhanced EAs. These efforts highlight the functional parallels between Transformer components and EA operations, confirming our conceptual analogy and demonstrating the potential for progress through idea sharing between advanced Transformers and reproduction.
@@ 1.4 Position embedding and selection
The attention matrix is influenced by both token and position relations. The selection matrix is usually designed based on fitness relations. In complex fitness landscapes, additional considerations such as genetic similarity among individuals are factored into the selection matrix. For instance, in multi-modal optimization with multiple global optima [67, 68], individual distances within the search space are employed as a criterion to preserve population diversity during selection. This ensures that the population does not converge prematurely to a single optimum. The selection, in this context, incorporates both individual relations (e.g., distances between individuals) and fitness relations (e.g., fitness rank). Essentially, the similarity between attention and selection matrices in handling relational information stems from the analogy drawn to tokens and individuals, as well as to positions and fitness. Existing Transformer-enhanced EA merges individual embeddings with fitness embeddings, echoing how token embeddings are combined with positional embeddings [31]. This practice serves as a support for our conceptual analogy. Advanced attention mechanisms, such as those incorporating rotational positional embedding [39], have the potential to enhance the performance of this operation in Transformer-enhanced EAs.
@@ 1.5 Model training and parameter adaptation
where $p_{i:N}$ refers to the $i$th ranked individual in a population with $N$ individuals based on fitness. The first term is the mean update, while the second term is the ranking-$N$ update of the covariance matrix. Recently, Transformer-enhanced EAs adaptively updated parameters from optimization experiences on a set of optimization tasks, improving the generalization ability of EAs on new tasks. Common methods include pre-training [31, 33] and meta-learning [42, 43]. In pre-training, optimization experiences for multi-objective optimization [31] consist of the population and their fitness generated by multi-objective EAs on numerous benchmarks. Optimization experiences for GP [33] include incremental changes in files submitted by humans to version control systems like Github. Meta-learning [42, 43] adaptively updates parameters by optimizing the average performance of EAs across a set of tasks. Regrettably, no comprehensive study has compared these two methods within a single framework. Furthermore, well-trained LLMs are directly utilized as reproduction operators with human-like experience [33, 71]. Prompts based on historical populations are constructed to guide LLMs in generating the desired output population [37, 33].
@@
Despite differing implementation strategies, LLMs and EAs converge on a shared fundamental objective: revealing the underlying probability distributions within data, thereby facilitating the learning and exploration of knowledge. In pre-training and supervised fine-tuning, LLMs construct conditional probability distributions through the accurate prediction of tokens, learning vast pre-existing knowledge. Reinforcement learning adjusts the LLM parameters based on the rewards. Evolutionary fine-tuning automatically searches for high-quality prompts or configurations to improve output quality [15, 72]. These two paradigms explore new knowledge specific to the target task. Learning and exploration jointly ensure the generative and generalization abilities of LLMs. EAs shape probability distributions based on fitness evaluated via real-time individual-environment interaction. In Transformer/LLM-enhanced EAs, models learn from existing optimization experiences or human-like experiences, endowing EAs with powerful learning capabilities. Additionally, LLM training works in the parameter space, while evolutionary fine-tuning extends to the language space. EAs operate both in the search space (e.g., GA) and the parameter space (e.g., CMAES). The aforementioned analogy provides a reasonable motivation for interdisciplinary research: merging the exceptional learning capabilities of LLMs with the remarkable exploration abilities of EAs can foster advancements in their respective fields.
@@
In (12), the hyperparameter $\mu$ is carefully tuned manually, as it affects the generalization ability of LLMs. The loss trade-off is modeled as a multi-objective optimization problem [73, 74]. Applying advanced multi-objective EAs aids in creating stronger supervised fine-tuning paradigms. Recent studies [64] show that ES has advantages over gradient-based reinforcement learning for long episodes with very many time steps. ES is promising as an alternative to reinforcement learning for training LLMs in multi-turn dialogue systems.
@@
1.6 Summary
Figure 3: Conceptual parallels between large language models and evolutionary algorithms: inspiring novel ideas and technical advancements.
Despite their independent development, LLMs and EAs share certain conceptual parallels. The parallels have inspired novel ideas and technical advancements, as outlined in Fig. 3. From a macro perspective, the parallels between LLMs and EAs provide a conceptual framework that can inspire the development of artificial agents capable of learning from established knowledge while continuously exploring new knowledge. For instance, recent studies have demonstrated that integrating EAs with LLMs can significantly enhance their performance and expand their application scope [15, 35, 37, 41, 31, 33, 42, 43, 40]. However, a unified paradigm with one-to-one key feature correspondence has not emerged. We stress that the analogy aims not to validate these parallels mathematically but to provide researchers with a pathway to enhance current technical studies. In existing efforts to integrate EAs and LLMs, evolutionary fine-tuning in black-box scenarios and LLM-enhanced EAs are receiving increasing attention. Next, this paper provides a comprehensive review of them to identify key challenges.